OpenStudy (anonymous):

Verify the identity. cot^3t/csct = cost(csc^2t-1)

OpenStudy (anonymous):

First, turn everything into sin and cos!

OpenStudy (anonymous):

OK cot = cos/sin, so cot^3 = cos^3 / sin ^3 csc = 1/sin so 1/csc = sin Left side: cot^3t/csct =(cos^3t / sin^3t) * sint =cos^3t / sin^2t Right side: 1+cot^2 = csc^2 csc^2 -1 = cot^2 so, right side: cost(csc^2t-1) =cost(cot^2t) =cost( cos^2t/sin^2t) =cos^3t/sin^2t Both sides are identical!